# Safe-SAIL: Towards a Fine-grained Safety Landscape of Large Language Models via Sparse Autoencoder Interpretation Framework

### v2

@@ 3.1.1 Settings
We select TopKReLU as the activation function because it allows easy control of the sparsity levels through the hyperparameter $k$. In our experiments, we chose k=20, 200, 500, 2000.
@@
The expansion factor is fixed at 10, which is based on previous work in SAEBench(Karvonen et al. 2025), where the SAE was evaluated with expansion factors of 4, 16, and 32. For input signals with 2048 dimensions, an expansion factor of 10 is a reasonable choice.
@@
To identify neurons associated with safety concepts in Qwen2.5-3B-Instruct, we intentionally selected potentially risky texts from our routine business traffic during the synthesis of the training data. Explanation data, separated from training data, is constructed by 200k queries mixed of 25% risky content, 10% random white queries and 65% randomly from public dataset The Pile(Gao et al. 2020). Evaluation data is constructed using Concept Contrastive Query Pairs method, which consists of 10,000 pairs across four safety domains: politics, pornography, violence and terror.
@@
We evaluate SAEs with existing interpretability metrics including $k$-Sparse Probing (Gurnee et al. 2023) and 1d-Probe (Gao et al. 2024), comparing with our own metrics on evaluation dataset.
@@ 3.1.2 Results
In terms of interpretability metrics, our proposed indicators demonstrate consistent trends across various safety domains (Figure 3), aligning more closely with the variability in neuron counts. It can be observed in Figure 5 that the effectiveness of k-Sparse Probing is significantly influenced by the choice of $k$, and the top-k mechanism focuses solely on the top neurons’ contribution to semantic classification, which fails to capture the overall representation of SAE. Furthermore, the 1d-Probe’s calculation of minimum cross-entropy loss reveals considerable instability, heavily dependent on the data, necessitating a large number of categories to yield effective results.
We find that concept-specific interpretability, is characterized by a higher number of neurons and more detailed explanations. This suggests a divergence between the optimal sparsity for concept-specific interpretability and that for minimal feature interference. According to earlier studies (Gao et al. 2024) and also illustrated in Figure 6, the effect of feature interference diminishes as more features are included in the reconstruction of the signal, up to a point where the features become optimally orthogonal. Beyond this threshold, the effects of superposition begin to dominate(Ferrando et al. 2024). Importantly, the SAE achieves the best concept-specific interpretability at a sparser level than that needed for minimal feature interference.
Table 2: Statistics of neuron explanations based on different explainer models. The average correlation score (Avg CorrScore), derived from simulations, is reported alongside the proportion of neuron explanations with correlation scores exceeding 0.2 ($R_{corr>0.2}$).
Explainer Model Average CorrScore $R_{corr>0.2}$
@@
This is because concept-specific domains, such as safety domains, are small subspaces within the larger semantic space, where features typically span the subspaces of frequently occurring concepts. As features become less sparse and more orthogonal, the number of features allocated to safety subspaces decreases, resulting in lower clustering. This is reflected in fewer explained neurons and coarser granularity in the resulting explanations.
@@ 3.2 Explainer Model Selection
The explainer model plays a crucial role in analyzing neuron activation samples, drawing conclusions about activation patterns, and ultimately producing human-interpretable explanations. The selection of the explainer model directly impacts the quality of the neuron database.
@@ 3.2.1 Settings
We compare explanations of neurons derived from all quartile layers (0, 8, 17, 26, 35) generated by different LRM models: QwQ-32B(Qwen Team 2025), DeepSeek-R1(DeepSeek-AI et al. 2025), and Claude 3.7 sonnet(Anthropic 2025). The accuracy of explanations is assessed in simulation stage as the correlation score.
@@ 3.2.2 Results
Table 2 shows that DeepSeek-R1 output performs other models in terms of average correlation score and the percentage of correlation score exceeding 0.2. According to subsequent experiments in the simulation section, neuron behaviors represented by explanations above this threshold are deemed interpretable by humans. The correlation score in this experiment is within a reasonable range comparable to previous work (Lieberum et al. 2024). Surprisingly, when QwQ-32B is tasked with interpreting code data activation samples, its responses exhibit significant confusion, characterized by the repetition of meaningless phrases, garbled output, and random responses, ultimately hindering its effectiveness in completing the task.
@@ 3.3.1 Settings
In this section, we compare existing simulation methods. We conduct experiments on layer 17 of Qwen2.5-3B-Instruct, with 1058 safe-related neurons, and use QwQ-32B for simulation. For every neuron, we sample 20 data from each activation bin of activations, if available.
The methods we evaluate include: 1) All at once: present each token in a ‘token<tab>unknown’ format within a single prompt, and then examines the logits for the unknown tokens to calculate a predicted activation as the probabilities weighted sum over token 0 to 10; 2) Token-level simulation: present each token in a ‘token<tab>unknown’ format, but the predicted activation is directly obtained from the LRM’s output; 3) Segment-level simulation: the original query is split into several segments, and the LRM is queried to determine whether each segment is activated or not.
We also collect token-level human-labeled activations for randomly selected 200 neurons, which serve as the ground truth for simulation results. Metrics we use are the correlation coefficient (Pearson $r$ and Kendall $\tau$) with human-labeled CorrScore. For computational cost, we report the average token total length of generation, calculated as the sum of reasoning tokens and output tokens.
Figure 7: Correlations between different methods and human-labeled results (top row), correlations between Segment-level and Token-level simulation (bottom left), and computational cost by generated token number (bottom right). Compared to token-level simulation, our method could reduce resource usage by 55% while maintaining decent performance ($r=0.8$).
Figure 8: Simulation performance and efficiency for different segment numbers. The left figure shows Pearson’s $r$ compared to token-level simulation, and the right displays the mean number of generated tokens. The orange dashed line represents the number of generated tokens generated by token-level simulation.
@@ 3.3.2 Results
Results compared with human-labeled token-level simulation are shown in top row of Figure 7. The bottom left figure shows a strong correlation ($r=0.8$) between segment-level and token-level CorrScore. Although Segment-level simulation is a simplification of Token-level simulation, it still preserves considerable performance while reducing computational cost by roughly 55%. We also report the simulation performance and efficiency for different numbers of segments in our segment-level simulation methods in Figure 8, as an approximation of token-level simulation.
@@
4. Insights
We present exploratory analyses demonstrating the utility of our SAE-based neuron interpretation database in uncovering the internal representation of safety-critical concepts in large language models. Focusing on pornography as a representative harmful category, we investigate how models encode specific real-world entities, such as known pornographic websites. These findings validate the framework’s effectiveness and enable concept-driven, semantically meaningful auditing of model knowledge.
@@ 4.1 Activation Patterns Imply Knowledge
We investigate whether LLMs encode latent knowledge of adult websites through SAE neuron activation patterns. Using the prompt “What is the main function of {web_url}?”, we evaluated 20 verified pornographic URLs, recording neuron activations at the first few tokens after URL input across 50 inference runs per URL. Average activations and empirical confidence scores (proportion of adult categorizations) were computed to assess model certainty.
Our results (Figure 9) reveal strong links between neuron activations and model behavior. Three neurons align closely with adult content detection: 26_18429 responds to semantic content (e.g., explicit URLs with 100% confidence), while 17_1579 and 17_4828 track syntactic patterns (e.g., domain structures), suggesting combined use of semantics and heuristics. Notably, onlyfans.com deviates—despite high confidence, these neurons show minimal activation. This suggests either (1) reliance on other, unobserved features, or (2) weak internal association between OnlyFans and explicit adult content, revealing limitations in the model’s conceptual representation.
The findings reveal that 2-3 specific neurons capture critical aspects of the model’s decision-making process, with distinct roles in semantic vs. syntactic processing. Such neuronal signatures provide interpretable markers for understanding model cognition and predicting outputs in safety-related tasks.
@@
Figure 10: Differences in the neuron activation chains between an English prompt (a) and a Hindi prompt (b), revealing how internal model mechanisms contribute to language-specific safety vulnerabilities.
@@ 4.2 Model Inference Trajectories
Large language models are often seen as black boxes, but our cross-layer neuron database enables fine-grained analysis of their internal representations. By tracing neuron activations across layers, we observe a clear progression: from local feature detection (e.g., keyword recognition) to structured reasoning (e.g., integrating semantics and context).
@@
In a case study on child sexual abuse related input (Figure 10), we identify a coherent processing pipeline: word-level detection (e.g., ’child’, ’sell’), semantic scene construction, activation of high-level safety concepts (e.g. transaction, sexual exploitation), and finally a safe refusal. The alignment between neuron semantics and model behavior shows that safety responses emerge from an interpretable, concept-driven reasoning chain—rather than arbitrary outputs. Furthermore, by examining neuron activation patterns across different languages, we gain interesting insights into the underlying mechanisms that give rise to safety vulnerabilities when the model processes low-resource languages. As shown in Figure 10, malicious Hindi inputs fail to trigger safe responses because the model lacks understanding of concepts such as child sexual abuse material and sexual exploitation-evident in the weak or absent activation of relevant neurons within the activation chain.
5. Conclusion
@@
We introduce a novel SAE interpretation framework that not only generates more granular safety neuron explanations but also reduces explanation costs by half. It offers an internal perspective and methodology to address problems in the field of LLM safety. Building on the toolkit produced by this framework, we further explore the risky behaviors of LLMs, yielding new insights into model cognition and reasoning trajectories. These findings enrich our understanding of LLM safety and lay the basis for future research in this area. We hope that by providing the toolkit including SAE checkpoints and a safety-tagged neuron database, our work will inspire greater interest in the field of LLM safety among researchers and equip scholars with new analytical tools.
Appendix A Appendix
A. Related Work
Sparse autoencoders
@@ A. Related Work
Sparse autoencoders are designed to transform an input signal, typically taken from MLP output or Residual Stream output, into a higher-dimensional representation. After a non-linear activation function, the encoded features are decoded back to reconstruct the input. Previous work on Sparse Autoencoders (SAEs) has explored various approaches to balance reconstruction accuracy and feature sparsity(Rajamanoharan et al. 2024; Gao et al. 2024; Bussmann, Leask, and Nanda 2024; Karvonen et al. 2024; Bussmann et al. 2025). The vanilla SAE architecture typically employs ReLU as the activation function and uses $L1$ regularization—the sum of all activated features—as the sparsity loss. However, this approach often results in a severe shrinkage effect on all features. Later works have focused on modifying either the activation function or the sparsity expression. TopKReLU(Gao et al. 2024) alters the activation function by selecting only the top-k features for signal reconstruction, making the sparsity level fixed. JumpReLU(Rajamanoharan et al. 2024) divides the activation function into two gated routes and penalizes only the binarized results on one route while preserving the feature magnitude on the other. While these methods focus on minimizing reconstruction loss at a certain level of sparsity, they seldom investigate the ultimate effect that reconstruction quality and feature sparsity have on the actual interpretability of the learned features.
LLM scopes
@@
Recent studies have expanded the application of Sparse Autoencoders (SAEs) to various layers of large language models, providing comprehensive insights into their internal representations. For instance, GemmaScope(Lieberum et al. 2024) applied SAE training to the Attention, MLP, and Residual layers of both Gemma-2B and Gemma-9B models. This approach allowed for a detailed examination of feature representations across different model components and scales. Similarly, LlamaScope(He et al. 2024) extended this methodology to the entire layer structure of the Llama3.1-8B-Instruct model, offering a holistic view of the model’s internal mechanisms. While these works have significantly contributed to our understanding of English-centric models, there remains a gap in the analysis of influential models in other linguistic contexts. Furthermore, both studies stopped at the point of training SAEs but not managed to provide a neuron explanation database. Our work addresses this gap by focusing on Qwen2.5-3B-Instruct, a prominent model in the Chinese language domain. By applying SAE training and safety-alignment related neuron explanation to this model, we aim to provide valuable insights into the internal representations of the Qwen series, which has had a substantial impact in Chinese natural language processing tasks.
Interpretation pipelines

### v3

@@ 3.1.1 Settings
We select TopKReLU as the activation function because it allows easy control of the sparsity levels through the hyperparameter $k$. In our experiments, we chose kk=20, 200, 500, 2000.
@@
We fix the expansion factor to 10, motivated by insights from Karvonen et al. (2025), who evaluate expansion factors across input dimensionalities. Given our 2048-dimensional inputs, this setting provides a balanced and appropriate configuration.
@@
Threshold Selection
For the $L_{0,t}$ metric, we employ a threshold of $t=0.25$. This value was empirically selected as it effectively enables distinguishability between all SAE configurations compared in our study.
@@
Data to train SAEs comprises query-response pairs covering politics, pornography, violence, and terrorism. Explanation data, separated from training data, is constructed by 200k queries mixed of 25% risky content, 10% random not risk-related queries and 65% randomly from public dataset The Pile Gao et al. (2020). Evaluation data is constructed using Concept Contrastive Query Pairs, which consists of 10,000 pairs across four safety domains: politics, pornography, violence and terror.
@@
We evaluate SAEs with existing interpretability metrics including $k$-Sparse Probing Gurnee et al. (2023) and 1d-Probe Gao et al. (2024), comparing with our own metrics on evaluation dataset.
@@ 3.1.2 Results
In terms of interpretability metrics, our proposed indicators demonstrate consistent trends across various safety domains (Figure 3), aligning more closely with the variability in feature counts. It can be observed in Figure 5 that the effectiveness of k-Sparse Probing is significantly influenced by the choice of $k$, and the top-k mechanism focuses solely on the top features’ contribution to semantic classification, which fails to capture the overall representation of SAE. Furthermore, the 1d-Probe’s calculation of minimum cross-entropy loss reveals considerable instability, heavily dependent on the data, necessitating a large number of categories to yield effective results.
We find that domain-specific interpretability, is characterized by a higher number of features and more detailed explanations. This suggests a divergence between the optimal sparsity for domain-specific interpretability and that for minimal feature interference. According to earlier studies (Gao et al., 2024) and also illustrated in Figure 6, the effect of feature interference diminishes as more features are included in the reconstruction of the signal, up to a point where the features become optimally orthogonal. Beyond this threshold, the effects of superposition begin to dominate Ferrando et al. (2024). Importantly, the SAE achieves the best domain-specific interpretability at a sparser level than that needed for minimal feature interference. This is because safety domains are small subspaces within the larger semantic space, where features typically span the subspaces of frequently occurring concepts. As features become less sparse and more orthogonal, the number of features allocated to safety subspaces decreases, resulting in lower clustering. This is reflected in fewer explained features and coarser granularity in the resulting explanations.
Explainer Model Avg. CorrScore $R_{corr>0.2}$
@@
Table 2: Statistics of feature explanations based on different explainer models. Avg. CorrScore is the average correlation score derived from simulations, $R_{corr>0.2}$ is the proportion of feature explanations with correlation scores exceeding 0.2.
@@ 3.2.1 Settings
We compare explanations of features derived from all quartile layers (0, 8, 17, 26, 35) generated by different LRM models: QwQ-32B Qwen Team (2025), DeepSeek-R1 DeepSeek-AI et al. (2025), and Claude 3.7 sonnet Anthropic (2025). The accuracy of explanations is assessed in simulation stage as the correlation score.
@@ 3.2.2 Results
Table 2 shows that DeepSeek-R1 outperforms other models in terms of average correlation score and the percentage of correlation score exceeding 0.2. According to subsequent experiments in the simulation section, feature behaviors represented by explanations above this threshold are deemed interpretable by humans. The correlation score in this experiment is within a reasonable range comparable to previous work Lieberum et al. (2024). Surprisingly, when QwQ-32B is tasked with interpreting code data activation samples, its responses exhibit significant confusion, characterized by the repetition of meaningless phrases, garbled output, and random responses.
@@ 3.3.1 Settings
In this section we compare existing simulation methods. We conduct experiments on layer 17 of Qwen2.5-3B-Instruct, with 1058 safe-related features, and use QwQ-32B for simulation. For every feature, we sample 20 data from each activation bin of activations, if available.
The methods we evaluate include: 1) All at once: present each token in a ‘token<tab>unknown’ format within a single prompt, and then examines the logits for the unknown tokens to calculate a predicted activation as the probabilities weighted sum over token 0 to 10; 2) Token-level simulation: present each token in a ‘token<tab>unknown’ format, but the predicted activation is directly obtained from the LRM’s output; 3) Segment-level simulation: the original query is split into $n$ segments, and the LRM is instructed to determine whether each segment is activated or not. For the remainder of this section, we refer to these two primary methods as TLS and SLS, respectively.
We also collect token-level human-labeled activations for randomly selected 200 features, which serve as the ground truth for simulation results. Metrics we use are the correlation coefficient (Pearson $r$ and Kendall $\tau$) with human-labeled CorrScore. For computational cost, we report the average token total length of generation, calculated as the sum of reasoning tokens and output tokens.
Figure 7: Correlations between different methods and human-labeled results (top row), correlations between SLS and TLS (bottom left), and computational cost by generated token number (bottom right). Compared to TLS, our method could reduce resource usage by 55% while maintaining decent performance ($r=0.8$).
Figure 8: Simulation performance and efficiency for different segment numbers. The left figure shows Pearson’s $r$ compared to TLS, and the right displays the mean number of generated tokens. The orange dashed line represents the number of generated tokens generated by TLS.
@@ 3.3.2 Results
Results compared with human-labeled TLS are shown in top row of Figure 7. The bottom left figure shows a strong correlation ($r=0.8$) between segment-level and token-level CorrScore when we choose $n=8$. Although SLS is a simplification of TLS, it still preserves considerable performance while reducing computational cost by roughly 55%. We also report the simulation performance and efficiency for different numbers of segments in SLS in Figure 8, as an approximation of TLS.
@@
4 Insights
In this section, we present exploratory analyses on our SAE-based feature interpretation database to uncover safety-critical concepts.
@@ 4.1 Activation Patterns Imply Knowledge
We take pornography as a representative harmful category and investigate how models encode specific real-world entities, such as known pornographic websites, through SAE feature activation patterns. Using the prompt “What is the main function of {web_url}?”, we evaluate 20 pornographic URLs, recording feature activations at the first few tokens after URL input across 50 inference runs per URL. Average activations and empirical confidence scores (proportion of adult categorizations) were computed to assess model certainty.
Our results (Figure 9) reveal strong links between feature activations and model behavior. Three features align closely with adult content detection: 26_18429 responds to semantic content (e.g., explicit URLs with 100% confidence), while 17_1579 and 17_4828 track syntactic patterns (e.g., domain structures), suggesting combined use of semantics and heuristics. Notably, onlyfans.com deviates—despite high confidence, these features show minimal activation. This suggests either (1) reliance on other, unobserved features, or (2) weak internal association between OnlyFans and explicit adult content. The findings reveal that 2-3 specific features capture critical aspects of the model’s decision-making process, with distinct roles in semantic v.s. syntactic processing. Such feature signatures provide interpretable markers for understanding model cognition and predicting outputs in safety-related tasks. See more details on related feature explanation and additional results in Appendix E.
@@
Figure 10: Differences in feature activation chains between an English prompt (a) and a Hindi prompt (b), highlighting how language-specific processing pathways may contribute to distinct safety vulnerabilities.
@@ 4.2 Model Inference Trajectories
Our cross-layer feature database enables fine-grained analysis of LLM internal representations. By tracing feature activations across layers, we observe a clear progression: from local feature detection (e.g., keyword recognition) to structured reasoning (e.g., integrating semantics and context).
@@
In a case study on child sexual abuse related input (Figure 10(a)), we identify a coherent processing pipeline: word-level detection (e.g., ‘child’, ‘sell’), semantic scene construction, activation of high-level safety concepts (e.g., transaction, sexual exploitation), and finally a safe refusal. The alignment between feature semantics and model behavior shows that safety responses emerge from an interpretable, concept-driven reasoning chain—rather than arbitrary outputs. Furthermore, by examining feature activation patterns across different languages, we gain interesting insights into the underlying mechanisms that give rise to safety vulnerabilities when the model processes low-resource languages. As shown in Figure 10(b), malicious Hindi inputs fail to trigger safe responses because the model lacks understanding of concepts such as child sexual abuse material and sexual exploitation-evident in the weak or absent activation of relevant features within the activation chain. See experiments on other languages in Appendix F.
5 Conclusion
@@
We introduce a novel SAE interpretation framework that not only generates more granular safety feature explanations but also reduces explanation costs by half. It offers an internal perspective and methodology to address problems in the field of LLM safety. Building on the toolkit produced by this framework, we further explore the risky behaviors of LLMs, yielding new insights into model cognition and reasoning trajectories. We hope that by providing the SAE checkpoints and a safety-tagged feature database, our work will inspire greater interest in the field of LLM safety with new analytical tools.
Limitations
Safe-SAIL’s current scope presents several opportunities for extension. While our safety feature database covers four major risk domains, extending it to emerging safety concerns such as privacy leakage, model self-replication, or deceptive alignment would broaden its applicability. Additionally, explanation quality depends on the explainer model’s capabilities, which may vary across languages and technical domains. Our segment-level simulation offers substantial efficiency gains but approximates token-level dynamics; applications requiring maximal granularity may benefit from hybrid approaches. Similarly, evaluation metrics rely on empirically validated thresholds that could benefit from automated domain-adaptive tuning. Finally, while our analyses primarily identify strong correlational patterns, establishing causal mechanisms through controlled interventions remains an important next step. We conduct preliminary neuron-level steering experiments on a subset of features and find that intervening on several identified neurons can measurably influence model behavior in directions consistent with their semantic interpretations (Appendix G). These results provide partial support for the causal relevance of some features, but they remain limited in scope and do not yet constitute a systematic causal validation of the broader feature set. We view these not as fundamental constraints but as natural avenues for scaling this interpretability paradigm.
Ethical Considerations
This work involves the analysis of safety-critical behaviors in large language models , including exposure to and interpretation of offensive, harmful, and illegal content such as pornography, violence, terrorism, and human trafficking. We acknowledge the ethical sensitivity of this research and have taken multiple measures to mitigate potential harms.
We handle sensitive content responsibly: only minimal, necessary examples appear in the main text. Our released toolkit includes only semantic interpretations, not raw toxic data, to limit misuse. The framework is designed solely to improve LLM safety through mechanistic interpretability, and all open-sourced components (SAE checkpoints, tools, etc.) are intended for defensive research. We mitigate potential misuse by avoiding release of jailbreaking-enabling assets, focusing on safety insights (e.g., multilingual failure modes), and restricting access in our demo. Finally, our multilingual findings aim to promote inclusive safety training—not discriminatory practices.
Due to data privacy and safety issues, our dataset used to train SAEs will not be made public. As for the use of AI assistants, beyond the AI models explicitly described in the experiments, we only used AI tools for language polishing to improve the clarity and fluency of the manuscript. All research ideas, experimental design, and analysis were conducted independently by the authors.
We conducted manual annotation as part of our experiments using our company’s in-house annotation platform. The annotators were professional third-party contractors employed by the company, compensated at a monthly rate of \$1,800, which is comparable to local jurisdictions. The annotation tasks involved no personal or private data.
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (62502435) and the Zhejiang Provincial Natural Science Foundation (LQN26F020002).
We would like to thank the team members from Alibaba Group for their contributions to the development of the Diagnose Toolkit and the companion product. In particular, we thank Wenchao Yang and Sha Xu for product support; Wendi Jia and Hongbin Li for frontend development; Song Liu and Hongwei Wu for backend development; Ke Zhang and Yushi Ma for design support; and Kun Huang for testing. We also extend special thanks to Boxuan Wang from Zhejiang University for his support and contributions throughout the project.
Appendix A Related Work
A.1 Sparse autoencoders
@@
Sparse autoencoders (SAEs) Bricken et al. (2023); Cunningham et al. (2023) are designed to transform an input signal, typically taken from MLP output or residual stream output, into a higher-dimensional representation; after a non-linear activation function, the encoded features are decoded back to reconstruct the input. Previous work on SAEs has explored various approaches to balance reconstruction accuracy and feature sparsity Rajamanoharan et al. (2024); Gao et al. (2024); Bussmann et al. (2024); Karvonen et al. (2024); Bussmann et al. (2025). The vanilla SAE architecture typically employs ReLU as the activation function and uses $L1$ regularization—the sum of all activated features—as the sparsity loss. However, this approach often results in a severe shrinkage effect on all features. Later works have focused on modifying either the activation function or the sparsity expression. TopKReLUGao et al. (2024) alters the activation function by selecting only the top-k features for signal reconstruction, making the sparsity level fixed. JumpReLURajamanoharan et al. (2024) divides the activation function into two gated routes and penalizes only the binarized results on one route while preserving the feature magnitude on the other. While these methods focus on minimizing reconstruction loss at a certain level of sparsity, they have not investigated the ultimate effect that reconstruction quality and feature sparsity have on the actual interpretability of the learned features.
A.2 LLM scopes
@@
Recent studies have expanded the application of SAEs to various layers of large language models, providing comprehensive insights into their internal representations. For instance, GemmaScope Lieberum et al. (2024) applied SAE training to the Attention, MLP, and residual layers of both Gemma-2B and Gemma-9B models. Similarly, LlamaScope He et al. (2024) extended this methodology to the entire layer structure of the Llama3.1-8B-Instruct model. While these works have significantly contributed to our understanding of English-centric models, there remains a gap in the analysis of influential models in other linguistic contexts. Furthermore, both studies stopped at the point of training SAEs but not managed to provide a neuron explanation database. Our work builds on Qwen2.5-3B-Instruct by applying SAE training and safety-alignment related neuron explanation to this model.
A.3 Interpretation pipelines